### v1

@@ 4.3.5 Multi-epoch radio transient surveys (2010s)
The 2010s saw the emergence of designed-for-purpose multi-epoch radio transient surveys. Some of the first of these were conducted with the Allen Telescope Array (ATA; deboer_allen_2006), an instrument designed for SETI searches. croft_allen_2011 published results from the 12-epoch carried out a 12-epoch survey, 690 deg2 survey (ATATS) at 1.4 GHz. bower_search_2011 and croft_allen_2013 presented results from the PiGSS survey at 3 GHz. Neither of these surveys detected any transients (which they defined as varying by a factor of >10>10). However, they set more stringent limits on transient rates. Unfortunately, due to funding changes, radio imaging surveys did not continue after these initial results, and the full potential of the array as a transient machine was not realised.
@@ 4.3.8 High frequency surveys
All of the surveys discussed so far have been conducted at or below frequencies of a few gigahertz. At higher radio frequencies (and in the millimetre regime), telescope fields of view are much smaller. For example, the field of view of the Australia Telescope Compact Array is $\sim 2$ arcminutes at 20 GHz, and the field of view of the Atacama Large Millimeter Array (ALMA; wootten_atacama_2009) is $\sim 19\hbox{${}^{\prime\prime}$}$ at 300 GHz. As a result, there are very few large scale continuum surveys at frequencies >10>10 GHz. Two examples are the 9C survey (waldram_9c_2003) covering 520 deg2 at 15 GHz and the AT20G survey (murphy_australia_2010) covering 20 626 deg2 at 20 GHz. Two conclusions from these projects were that (i) the overall properties of the source population at high frequencies can not be extrapolated from the $\sim 1$ GHz population; (ii) the high frequency radio sky has a comparable level of variability to gigahertz frequencies: ($\sim 5\%$ of bright sources were variable on 121-2 year timescales sadler_properties_2006).
@@ 4.3.9 Galactic Centre and Galactic plane surveys
The next major long-term Galactic centre monitoring program was conducted by hyman_low-frequency_2002. This project used archival and new widefield images of the Galactic centre taken with the VLA at 0.33 GHz, with targeted follow-up of objects of interest at 1.4 and 4.8 GHz. They introduced the nomenclature Galactic Centre Radio Transient (GCRT) to describe these sources, as their nature was unknown. Subsequent work by hyman_powerful_2005; hyman_new_2006; hyman_faint_2007; hyman_gcrt_2009 resulted in the discovery of three GRCTs. These sources shared some common properties (highly variable, steep spectrum, circularly polarised) but their identity has remained elusive (see Section 3.9).
@@ 4.4 Current large-scale radio transient surveys
N repeats 1f ongoing ongoing 21e 120e 3 1 1 3.5
@@ 4.4.2 The MWA Galactic Plane Monitor
The main scientific goal of the survey is to discovery long-period transients. The survey parameters are briefly described in the Methods section of hurley-walker_long-period_2023, which reports the discovery of GPM J1839-10, a new long period transient, and will be fully described in a paper in preparation.
@@ 5.4 Short-timescale (‘fast imaging’) approaches
The current set of interferometers, with a large number of dishes or antennas, and hence much improved instantaneous $uv$-coverage have made imaging (or re-imaging) on short timescales possible. One example of this approach is the fast imaging of MeerKAT data on 8 s timescales by fijma_new_2024. Another example is the VASTER fast imaging pipeline (lee_searching_2025) which images ASKAP data on 10 s and 10 min timescales. Both of these projects ran the fast imaging in offline, batch-mode.
@@ 5.7.1 Radio lightcurve classification
Along similar lines, stewart_optical_2018 investigated radio versus optical flux density as a means of making a preliminary classification of radio transients. They showed that different source types do fall (broadly) into different regions of this two-parameter space. For example, quasars tend to have quite a high radio to optical flux density ratio, whereas stars typically have a much lower one. wang_discovery_2022 took this further and used radio to near-infrared (to be more robust against extinction) flux radio versus circular polarisation fraction, showing that the locations of pulsars and stellar sources (the two main classes of sources with significant polarisation) were very different. And wang_detection_2025 compared X-ray and radio fluxes (drawing on the radio/X-ray correlation plots for accreting sources; e.g. gallo_universal_2003; russell_reproducible_2016; ridder_radio_2023), showing that LPTs occupied an extreme region of that space distinct from other potentially related populations like low-mass X-ray binaries, stars, and cataclysmic variables, but did resemble magnetars. As with the timescale of variability, these metrics alone are not enough to definitively classify objects, but would be useful parts of a classification pipeline.
@@ 5.7.3 Citizen science projects
Citizen science projects (one of the first, and most influential of which, was SETI@HOME; werthimer_berkeley_2001) are designed to leverage the enthusiasm of members of the public to help complete large scale scientific data collection and classification, either using only their personal computer facilities (as in the case of SETI@HOME) or using the citizens themselves to help analyse data. In astronomy, a pioneering project in this space was Galaxy Zoo (lintott_galaxy_2008) which invited the general public to visually inspect and classify the morphology of nearly million galaxies from the Sloan Digital Sky survey (lintott_galaxy_2011). The success of the initial Galaxy Zoo project led it to broaden its remit to the Zooniverse. This platform, and other similar approaches, have been applied to a wide range of classification tasks including gravitational wave glitch identification (zevin_gravity_2017), exoplanet discovery (christiansen_k2-138_2018), and radio galaxy classification (lukic_radio_2018).
@@ 6.1.4 $\eta$-V parameters
Figure 9: Standard $\eta$-$V$ plot for identifying variable sources, following swinbank_lofar_2015. The data here were simulated and consist entirely of Gaussian noise with constant amplitude. The distribution of source brightnesses are a power-law with cumulative distribution $N(>S)\propto S^{-1.5}$, appropriate for standard candles in a Euclidean universe (e.g., hoyle_counting_1961; longair_luminosity_1965), with SNR ranging from 6 to 2000. The points are coloured by SNR (with a colour-scale to the right). We also show marginal distributions for $\eta$ (top) and $V$ (right), along with theoretical predictions (solid red lines). Finally, we show the line $V_{\rm max}=\sqrt{\eta}/{\rm SNR}_{\rm min}$ which bounds the top of the distribution (red dashed line).
@@ 6.1.5 Other variability metrics
Perhaps the simplest is the temporal gradient $\nabla_{F}\equiv dS/dt$ used by bell_murchison_2019, who also divided by the gradient by its uncertainty, $\nabla_{S}\equiv\nabla_{F}/\sigma_{F}$ to identify statistically significant changes. They also found that higher-order polynomial changes were not needed to describe the data.
@@ 7.1 The SKA Observatory
The SKA Observatory (SKAO) consists of two telescopes (SKA-Low and SKA-Mid), currently under construction [^17] The specifications given in this Section are based on the official SKAO documents available in this Zenodo collection: https://zenodo.org/communities/skaokeydocuments/, in particular the ‘SKAO Staged Delivery, Array Assemblies And Layouts’ document: seethapuram_skao_2025 and the ‘Design Baseline Description’: dewdney_ska1_2022.. SKA-Low will consist of 512 low frequency aperture array stations, spread over an area 80 km in diameter, located in Western Australia. Each station will have 256 wire antennas in a 38-m diameter area. It will observe in the frequency range 50–350 MHz. SKA-Mid will consist of 197 mid frequency dishes (64 from the existing MeerKAT array) spread over an area 180 km in diameter. It will be located in the Karoo region of South Africa. The telescope will operate in 6 frequency bands (3 initially) ranging from 580 MHz to 15.4 GHz. Full details and technical specifications of the telescopes are given in the Design Baseline Description and the SKAO Staged Delivery, Array Assemblies And Layouts documents and summarised in Table 6.
@@
Telescope LOFAR 2.0 LBAa SKA-Low LOFAR2.0 HBAa SKA-Midb DSA-2000 ngVLAc
@@
The science case for the SKAO telescopes has been outlined in two paper collections. Firstly, in ‘Science with the Square Kilometre Array’ (Carilli & Rawlings 2004) and then, a decade later, in ‘Advancing Astrophysics with the Square Kilometre Array’ (braun_advancing_2015). A new science book, stemming from the 2025 SKAO General Science Meeting, ‘A new era in astrophysics: Preparing for early science with the SKAO’ is currently in preparation.
The exact operating model of the SKAO telescopes (including how time is allocated, and potential survey design) is yet to be determined. However, the current plan is that PI-led science would start in 2030 on AA and larger observing programs led by the Key Science Projects would start several years later, on AA4.
@@ 7.3 The next-generation VLA
LOFAR 2.0202020See https://www.lofar.eu/lofar2-0-documentation/. is a major upgrade to the LOFAR telescope. Two new stations will be added. New low-noise amplifiers will be added to all antennas (wang_real-time_2025), and processing improvements will allow simultaneous use of both LBAs and HBAs simultaneously, or using two beams for each HBA station. There are also improvements in the clock distribution and central processing facility used for calibration and imaging, as well as a 24/7 transient buffer for imaging and localisation. The net result will be a considerably more powerful and flexible system capable of optimising its operation for many different transient cases (timescales, sensitivities, etc), as well as general improvements in image fidelity and processing (enabling faster followup with better localisation).
@@ 8 Summary and outlook
  • Multiwavelength coverage: Multiwavelength (optical, near-infrared, X-ray, gamma-ray) survey data is vital for assessing the nature of new discoveries. In general this does not have to be contemporaneous, although this is beneficial for some scientific goals. For example, stellar flares show significant strucure with an event, so modelling requires multi-wavelength data, but triggering is impractical. For synchrotron transients, optical/IR data is required to identify a host galaxy, and archival high-energy data can be vital to determining whether it is a novel event like an orphan afterglow.
@@ Acknowledgements
We had useful interactions and discussions with many people in the preparation of this review. Particular thanks to Gemma Anderson, Shari Breen, Manisha Caleb, Lauren Carter, Shami Chatterjee, Barnali Das, Paul Demorest, Iris de Ruiter, Dougal Dobie, Laura Driessen, Ashna Gulati, Gregg Hallinan, Natasha Hurley-Walker, Francois Kapp, Emil Lenc, Eric Murphy, Joshua Pritchard, Kovi Rose, Antonia Rowlinson, Sarrvesh Sridhar, Randall Wayth, Michael Wheatland and Andrew Zic.

### v2

@@ 4.3.5 Multi-epoch radio transient surveys (2010s)
The 2010s saw the emergence of designed-for-purpose multi-epoch radio transient surveys. Some of the first of these were conducted with the Allen Telescope Array (ATA; deboer_allen_2006), an instrument designed for SETI searches. croft_allen_2011 published results from the 12-epoch, 690 deg2 survey (ATATS) at 1.4 GHz. bower_search_2011 and croft_allen_2013 presented results from the PiGSS survey at 3 GHz. Neither of these surveys detected any transients (which they defined as varying by a factor of >10>10). However, they set more stringent limits on transient rates. Unfortunately, due to funding changes, radio imaging surveys did not continue after these initial results, and the full potential of the array as a transient machine was not realised.
@@ 4.3.8 High frequency surveys
All of the surveys discussed so far have been conducted at or below frequencies of a few gigahertz. At higher radio frequencies (and in the millimetre regime), telescope fields of view are much smaller. For example, the field of view of the Australia Telescope Compact Array is $\sim 2$ arcminutes at 20 GHz, and the field of view of the Atacama Large Millimeter Array (ALMA; wootten_atacama_2009) is $\sim 19\hbox{${}^{\prime\prime}$}$ at 300 GHz. As a result, there are very few large scale continuum surveys at frequencies >10>10 GHz. Two examples are the 9C survey (waldram_9c_2003) covering 520 deg2 at 15 GHz and the AT20G survey (murphy_australia_2010) covering 20 626 deg2 at 20 GHz. Two conclusions from these projects were that (i) the overall properties of the source population at high frequencies can not be extrapolated from the $\sim 1$ GHz population; (ii) the high frequency radio sky has a comparable level of variability to gigahertz frequencies: ($\sim 5\%$ of bright sources were variable on 121-2 year timescales; sadler_properties_2006).
@@ 4.3.9 Galactic Centre and Galactic plane surveys
The next major long-term Galactic centre monitoring program was conducted by hyman_low-frequency_2002. This project used archival and new widefield images of the Galactic centre taken with the VLA at 0.33 GHz, with targeted follow-up of objects of interest at 1.4 and 4.8 GHz. They introduced the nomenclature Galactic Centre Radio Transient (GCRT) to describe these sources, as their nature was unknown. Subsequent work by hyman_powerful_2005; hyman_new_2006; hyman_faint_2007; hyman_gcrt_2009 resulted in the discovery of three GCRTs. These sources shared some common properties (highly variable, steep spectrum, circularly polarised) but their identity has remained elusive (see Section 3.9).
@@ 4.4 Current large-scale radio transient surveys
N repeats 1e ongoing ongoing 21f 120f 3 1 1 3.5
@@ 4.4.2 The MWA Galactic Plane Monitor
The main scientific goal of the survey is to discover long-period transients. The survey parameters are briefly described in the Methods section of hurley-walker_long-period_2023, which reports the discovery of GPM J1839-10, a new long period transient, and will be fully described in a paper in preparation.
@@ 5.4 Short-timescale (‘fast imaging’) approaches
The current set of interferometers, with a large number of dishes or antennas, and hence much improved instantaneous $uv$-coverage have made imaging (or re-imaging) on short timescales possible. One example of this approach is the fast imaging of MeerKAT data on 8 s timescales by fijma_new_2024. Another example is the VASTER fast imaging pipeline (lee_searching_2026) which images ASKAP data on 10 s and 10 min timescales. Both of these projects ran the fast imaging in offline, batch-mode.
@@ 5.7.1 Radio lightcurve classification
Along similar lines, stewart_optical_2018 investigated radio versus optical flux density as a means of making a preliminary classification of radio transients. They showed that different source types do fall (broadly) into different regions of this two-parameter space. For example, quasars tend to have quite a high radio to optical flux density ratio, whereas stars typically have a much lower one. wang_discovery_2022 took this further and used radio to near-infrared (to be more robust against extinction) flux ratio versus circular polarisation fraction, showing that the locations of pulsars and stellar sources (the two main classes of sources with significant polarisation) were very different. And wang_detection_2025 compared X-ray and radio fluxes (drawing on the radio/X-ray correlation plots for accreting sources; e.g. gallo_universal_2003; russell_reproducible_2016; ridder_radio_2023), showing that LPTs occupied an extreme region of that space distinct from other potentially related populations like low-mass X-ray binaries, stars, and cataclysmic variables, but did resemble magnetars. As with the timescale of variability, these metrics alone are not enough to definitively classify objects, but would be useful parts of a classification pipeline.
@@ 5.7.3 Citizen science projects
Citizen science projects (one of the first, and most influential of which, was SETI@HOME; werthimer_berkeley_2001) are designed to leverage the enthusiasm of members of the public to help complete large scale scientific data collection and classification, either using only their personal computer facilities (as in the case of SETI@HOME) or using the citizens themselves to help analyse data. In astronomy, a pioneering project in this space was Galaxy Zoo (lintott_galaxy_2008) which invited the general public to visually inspect and classify the morphology of nearly a million galaxies from the Sloan Digital Sky survey (lintott_galaxy_2011). The success of the initial Galaxy Zoo project led it to broaden its remit to the Zooniverse. This platform, and other similar approaches, have been applied to a wide range of classification tasks including gravitational wave glitch identification (zevin_gravity_2017), exoplanet discovery (christiansen_k2-138_2018), and radio galaxy classification (lukic_radio_2018).
@@ 6.1.4 $\eta$-V parameters
Figure 9: Standard $\eta$-$V$ plot for identifying variable sources, following swinbank_lofar_2015. The data here were simulated and consist entirely of Gaussian noise with constant amplitude. The distribution of source brightnesses are a power-law with cumulative distribution $N(>S)\propto S^{-1.5}$, appropriate for standard candles in a Euclidean universe (e.g., hoyle_counting_1961; longair_luminosity_1965), with SNR ranging from 6 to 2000. The points are coloured by SNR (with a colour-scale to the right). We also show marginal distributions for $\eta$ (top) and $V$ (right), along with theoretical predictions (solid orange lines). Finally, we show the line $V_{\rm max}=\sqrt{\eta}/{\rm SNR}_{\rm min}$ which bounds the top of the distribution (red dashed line).
@@ 6.1.5 Other variability metrics
Perhaps the simplest is the temporal gradient $\nabla_{F}\equiv dS/dt$ used by bell_murchison_2019, who also divided the gradient by its uncertainty, $\nabla_{S}\equiv\nabla_{F}/\sigma_{F}$ to identify statistically significant changes. They also found that higher-order polynomial changes were not needed to describe the data.
@@ 7.1 The SKA Observatory
The SKA Observatory (SKAO) consists of two telescopes (SKA-Low and SKA-Mid), currently under construction [^17] The specifications given in this Section are based on the official SKAO documents available in this Zenodo collection: https://zenodo.org/communities/skaokeydocuments/, in particular the ‘SKAO Staged Delivery, Array Assemblies And Layouts’ document: seethapuram_skao_2025 and the ‘Design Baseline Description’: dewdney_ska1_2022.. SKA-Low will consist of 512 low frequency aperture array stations, spread over an area 80 km in diameter, located in Western Australia. Each station will have 256 wire antennas in a 38-m diameter area. It will observe in the frequency range 50–350 MHz. SKA-Mid will consist of 197 mid frequency dishes (64 from the existing MeerKAT array) spread over an area 160 km in diameter. It will be located in the Karoo region of South Africa. The telescope will operate in 6 frequency bands (3 initially) ranging from 350 MHz to 15.4 GHz. Full details and technical specifications of the telescopes are given in the Design Baseline Description and the SKAO Staged Delivery, Array Assemblies And Layouts documents and summarised in Table 6.
@@
Telescope LOFAR 2.0 LBAa SKA Low LOFAR2.0 HBAa SKA Midb DSA-2000 ngVLAc
@@
The science case for the SKAO telescopes has been outlined in two paper collections. Firstly, in ‘Science with the Square Kilometre Array’ (carilli_science_2004), and then, a decade later, in ‘Advancing Astrophysics with the Square Kilometre Array’ (braun_advancing_2015). A new science book, stemming from the 2025 SKAO General Science Meeting, ‘A new era in astrophysics: Preparing for early science with the SKAO’ is currently in preparation.
The exact operating model of the SKAO telescopes (including how time is allocated, and potential survey design) is yet to be determined. However, the current plan is that PI-led science would start in 2030 on AA and larger observing programs led by the Key Science Projects would start several years later.
@@ 7.3 The next-generation VLA
LOFAR 2.0202020See https://www.lofar.eu/lofar2-0-documentation/. is a major upgrade to the LOFAR telescope. Two new stations will be added. New low-noise amplifiers will be added to all antennas (wang_real-time_2025), and processing improvements will allow simultaneous use of both LBAs and HBAs, or using two beams for each HBA station. There are also improvements in the clock distribution and central processing facility used for calibration and imaging, as well as a 24/7 transient buffer for imaging and localisation. The net result will be a considerably more powerful and flexible system capable of optimising its operation for many different transient cases (timescales, sensitivities, etc), as well as general improvements in image fidelity and processing (enabling faster followup with better localisation).
@@ 8 Summary and outlook
  • Multiwavelength coverage: Multiwavelength (optical, near-infrared, X-ray, gamma-ray) survey data is vital for assessing the nature of new discoveries. In general this does not have to be contemporaneous, although this is beneficial for some scientific goals. For example, stellar flares show significant structure within an event, so modelling requires multi-wavelength data, but triggering is impractical. For synchrotron transients, optical/IR data is required to identify a host galaxy, and archival high-energy data can be vital to determining whether it is a novel event like an orphan afterglow.
@@ Acknowledgements
We had useful interactions and discussions with many people in the preparation of this review. Particular thanks to Gemma Anderson, Shari Breen, Manisha Caleb, Fernando Camilo, Lauren Carter, Shami Chatterjee, Barnali Das, Paul Demorest, Iris de Ruiter, Dougal Dobie, Laura Driessen, Ashna Gulati, Gregg Hallinan, Natasha Hurley-Walker, Scott Hyman, Francois Kapp, Emil Lenc, Walter Max-Moerbeck, Eric Murphy, Joshua Pritchard, Kovi Rose, Antonia Rowlinson, Sarrvesh Sridhar, Randall Wayth, Michael Wheatland and Andrew Zic.